OpenStudy (anonymous):

If I buy 2 apples and 6 oranges, I will spend $5.60. However, if I buy 6 apples and 2 oranges, I will spend $4.80. How much does one orange cost? Write the answer like this example: 48cents.

OpenStudy (anonymous):

2a+6o=5.6 6a+2o=4.8 SOLVE FOR a and o

OpenStudy (anonymous):

2a+6o=5.6 6a+2o=4.8 SOLVE FOR a and o

OpenStudy (anonymous):

h

OpenStudy (anonymous):

huh ?

OpenStudy (anonymous):

can you solve a system of equations?

OpenStudy (anonymous):

No : (

OpenStudy (anonymous):

2a+6o=5.6 (1) 6a+2o=4.8 (2) multiply equation (1) by -3 -6a-18o=-16.8 6a+2o=4.8 add the 2 equations: -16o=-12 o=0.75$=75 cents